Circular No. 5791

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

NOVA AQUILAE 1993 T. Hirayama, National Astronomical Observatory, Tokyo, reports that

Minoru Yamamoto of Doi-cho, Okazaki-shi, Japan, has discovered an apparent nova on T-Max 400 film (+ PO0 green filter). The object is located at α = 19^{h}10^{m}.5, δ = +1^{o}28′ (equinox 1950.0). Magnitude estimates, visual unless otherwise noted: Apr. 17.780 UT, [10 (Yamamoto; photographic); 28.802, [12 (P. Camilleri, Cobram, Vic.; T-Max film); May 14.644 and 14.647, 7.6 (Yamamoto; photographic); 14.779, 7.6 (Y. Sakurai, Mito, Japan; Fuji HG-400 film); 15.009, 7.6 (M. Collins, Everton, Beds, U.K.; Kodak Tech Pan film); 15.604, 7.9 (Yamamoto); 15.648, 7.6 (Yamamoto; photographic); 16.541, 7.8 (Camilleri); 16.606, 8.0 (M. Yamada, Ishikawa, Japan); 16.640, 8.0 (Yamada); 16.653, 7.9 (M. Watanabe, Toyama, Japan); 17.19, 8.0 (M. Baldwin, Columbus, IN).

EX LUPI Observations by A. F. Jones, Nelson, New Zealand; and W. B. Al-

brecht, Pahala, HI, indicate that this irregular variable is now quite active. Visual estimates by Jones, using the revised sequence on IBVS 3808: 1993 Apr. 2.68 UT, 12.8; 14.73, 12.5; 16.46, 12.7; 17.73, 12.4; 18.68, 12.0; 19.64, 12.3; 20.42, 12.5; 21.68, 12.7; 24.68, 12.5; 25.46, 12.4; 27.49, 12.4; 28.69, 12.1; 29.38, 12.0; 29.73, 11.5; May 13.70, 11.4. Photometry by A. C. Gilmore and P. M. Kilmartin at Mt. John shows that the star is brighter and bluer than observed by Bastien and Mundt (1979, A.Ap. Suppl. 36,

57) in 1977 and 1979 and by Gilmore (IBVS 3808) on 1992 Oct. 17: 1993 May 1.54, V = 12.27, U–B = –0.50, B–V = +0.52, V–R = +0.54; 9.51, 12.81, –, +0.90, +0.71; 10.35, 12.69, –0.28, +0.83, –; 13.53, 11.58, –0.36, +0.39, +0.36.

SUPERNOVA 1993N IN UGC 5695 J. Mueller, Palomar Observatory, forwards the following report from

J. R. Mould: “SN 1993N is 7′′ west and 9′′ north of the galaxy nucleus. A spectrum obtained on May 14 UT by Mould and I. Smail with the Hale 5-m telescope (+ Double Spectrograph) shows Hα and Hβ in emission. These line profiles show narrow cores a few hundred km/s FWHM on broad pedestals 16 600 km/s FWZI for Hα. The Balmer lines are accompanied by very narrow [O III]. The galaxy redshift is 2940 km/s. ”

1993 May 17 Daniel W. E. Green